### v2

# CORE4D: A 4D Human-Object-Human Interaction Dataset for Collaborative Object REarrangement

Abstract
Understanding how humans cooperatively rearrange household objects is critical for VR/AR and human-robot interaction. However, in-depth studies on modeling these behaviors are under-researched due to the lack of relevant datasets. We fill this gap by presenting CORE4D, a novel large-scale 4D human-object-human interaction dataset focusing on collaborative object rearrangement, which encompasses diverse compositions of various object geometries, collaboration modes, and 3D scenes. With 1K human-object-human motion sequences captured in the real world, we enrich CORE4D by contributing an iterative collaboration retargeting strategy to augment motions to a variety of novel objects. Leveraging this approach, CORE4D comprises a total of 11K collaboration sequences spanning 3K real and virtual object shapes. Benefiting from extensive motion patterns provided by CORE4D, we benchmark two tasks aiming at generating human-object interaction: human-object motion forecasting and interaction synthesis. Extensive experiments demonstrate the effectiveness of our collaboration retargeting strategy and indicate that CORE4D has posed new challenges to existing human-object interaction generation methodologies.
CORE4D is a large-scale diverse human-object-human interaction dataset for collaborative object rearrangement, encompassing real-world and synthetic branches. CORE4D-Real captures 1K human-object-human mesh sequences with allocentric and egocentric visual signals, while CORE4D-Synthetic retargets real-world data onto 3K virtual object shapes formulating 10K motion sequences.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
Humans frequently rearrange household items through multi-person collaboration , such as moving a table or picking up an overturned chair together. Analyzing and synthesizing these diverse collaborative behaviors could be widely applicable in VR/AR, human-robot interaction [103, 73, 72], dexterous manipulation [11, 121, 104, 134] and humanoid manipulation [71, 20, 117, 50]. However, understanding and modeling these interactive motions have been limited due to the lack of large-scale, richly annotated datasets. Most existing human-object and hand-object interaction datasets focus on individual behaviors [92, 3, 57, 128, 24, 41, 135, 53, 59, 132] and two-person handovers [129, 109, 58]. But these datasets typically encompass a limited number of object instances, thus struggling to support generalizable interaction understanding across diverse object shapes. Scaling up precise human-object interaction data is challenging. While vision-based human-object motion tracking methods [114, 141, 115, 116] have advanced, they still struggle with low fidelity due to severe occlusions, which are common in multi-human collaborations. Additionally, mocap [41, 53] is expensive and hard to scale up to cover numerous objects involved in rearrangement. Our goal is to curate a large-scale category-level human-object-human (HOH) interaction dataset with high quality in a cost-efficient manner.
We observe that HOH collaborations mainly vary in two aspects: the temporal collaboration patterns of two humans and the spatial relations between human and object. The temporal collaboration patterns could vary widely depending on scene complexity, motion range, and collaboration mode. In contrast, the spatial relations between human and object tend to possess strong homogeneity when facing objects from the same category, e.g., two persons holding opposite sides of a chair. This allows for retargeting interactions involving one specific instance to another using automatic algorithms, avoiding the need to capture interactions with thousands of same-category objects in the real world. The above observations enable us to leverage expensive motion capture systems to capture only humans’ diverse temporal collaboration patterns, while relying on automatic spatial retargeting algorithms to enrich human-object spatial relations.
Using these insights, we build a novel large-scale dataset, CORE4D, encompassing a wide range of human-object interactions for collaborative object rearrangement. CORE4D includes various types of household objects, collaboration modes, and 3D environments. Our data acquisition strategy combines mocap-based capturing and synthetic retargeting, allowing us to scale the dataset effectively. The retargeting algorithm transfers spatial relation between human and object to novel object geometries while preserving temporal pattern of human collaboration. As a result, CORE4D includes 1K real-world motion sequences (CORE4D-Real) paired with videos and 3D scenes, as well as 10K synthetic collaboration sequences (CORE4D-Synthetic) covering 3K diverse object shapes.
We benchmark two tasks for generating human-object collaboration: (1) motion forecasting [14, 119] and (2) interaction synthesis [97, 53] on CORE4D, revealing challenges in modeling human behaviors, enhancing motion naturalness, and adapting to new object geometries. Ablation studies demonstrate the effectiveness of our hybrid data acquisition strategy, and the quality and value of CORE4D-Synthetic, highlighting its role in helping to improve existing motion generation methods. We further retarget interactions in CORE4D onto Unitree H1 [100] humanoid robot and use them to train humanoid box-lifting policies, showcasing the values of CORE4D in robot interaction skill learning.
In summary, our main contributions are threefold: (1) We present CORE4D, a large-scale 4D HOH interaction dataset for collaborative object rearrangement. (2) We propose a novel hybrid data acquisition method, incorporating real-world data capture and synthetic collaboration retargeting. (3) We benchmark two tasks for collaboration generation, revealing new challenges and research opportunities.
2 Related Work
2.1 Human-object Interaction Datasets
Tremendous progress has been made in constructing human-object interaction datasets. To study how humans interact with 3D scenes, various widely-used datasets record human movements and surrounding scenes separately, treating objects as static [4, 86, 32, 106, 33, 55, 18, 137, 139, 30, 107, 2, 38, 123, 19, 31, 95, 124, 43, 130, 42] or partially deformable [54] without pose changes. For dynamic objects, recent works [5, 13, 6, 45, 48, 98, 92, 17, 3, 39, 41, 135, 53, 109, 58, 140, 136, 56, 141, 63, 126] have captured human-object interaction behaviors with varying focuses. Table 1 generally summarizes the characteristics of 4D human-object-interaction datasets. To support research for vision-based human-object motion tracking and shape reconstruction, a line of datasets [17, 3, 39, 41, 135, 140, 136, 56] present human-object mesh annotations with multi-view RGB or RGBD signals. With the rapid development of human-robot cooperation, several works [5, 92, 109, 58] focus on specific action types such as grasping [92] and human-human handover [5, 109, 58]. Our dataset uniquely captures multi-person and object collaborative motions, category-level interactions, and both egocentric and allocentric views, offering comprehensive features with the inclusion of both real and synthetic datasets.
dataset multi- collaboration category- egocentric RGBD ##\##view mocap ##\##object ##\##sequence
human level
GRAB [92] - 57 -
GraviCap [17] 3 4 9
BEHAVE [3] 4 20 321
InterCap [39] 6 10 223
CHAIRS [41] 4 81 1.4K
HODome [135] 76 23 274
Li et.al. [53] - 15 6.1K
FORCE [140] 1 8 450
IMHD2 [141] 32 10 295
HIMO [63] - 53 3.4K
Carfi et.al. [5] 1 10 1.1K
HOH [109] 8 136 2.7K
CoChair [58] - 8 3.0K
HOI-M3 [136] 42 90 199
CORE4D-Real 5 37 1.0K
CORE4D-Synthetic - - 3.0K 10K
Table 1: Comparison of CORE4D with existing 4D human-object interaction datasets.
2.2 Human Interaction Retargeting
Human interaction retargeting focuses on applying human interactive motions to novel objects in human-object interaction scenarios. Existing methodologies [46, 85, 128, 88, 111, 9, 40, 115] are object-centric. They propose first finding contact correspondences between the source and the target objects and then adjusting human motion to touch specific regions on the target object via optimization. As crucial guidance for the result, contact correspondences are discovered by aligning either surface regions [85, 128, 111, 115], spatial maps [46, 40], distance fields [9], or neural descriptor fields [88] between the source and the target objects. These methods are all limited to objects with similar topology and scales. Our synthetic data generation strategy incorporates object-centric design [128] with novel human-centric contact selection, allowing adaptation to challenging objects using human priors.
2.3 Human-object Interaction Generation
Human-object interaction generation is an emerging research topic that aims to synthesize realistic human-object motions conditioned on surrounding 3D scenes, known object trajectories, or action types. To generate 3D human mesh snapshots interacting with scenes, POSA [34] and COINS [142] propose to leverage CVAE [89], while DreamHOI [145] provides an iterative NeRF [69] optimization approach. To further synthesize interactive motions, a line of works [91, 27, 112, 93, 139, 49, 138, 70] present auto-regressive manners [91, 139], diffusion models [49], or two-stage designs that first generates start and end poses and then interpolates motion in-between [27, 112, 93, 138]. Beyond static objects, a line of works further model object movements and generate integrated human-object interactions using diffusion models [119, 53] and GCN [122]. To generate human-object interactions under action descriptions, recent works [22, 52, 81, 110, 120, 113, 108, 131, 23, 42, 90] extract text features with pretrained CLIP encoders [83, 108, 131, 42, 90] or LLM planners [75, 99, 113, 23], using them to guide diffusion models [37].
3 Constructing CORE4D
CORE4D is a large-scale 4D human-object-human interaction dataset acquired in a novel hybrid scheme, comprising CORE4D-Real and CORE4D-Synthetic. CORE4D-Real is captured (Section 3.1) and annotated (Section 3.2) from authentic collaborative scenarios. As shown in Figure 2, it provides human-object-human poses, egocentric RGB videos, allocentric RGB-D videos, and 2D segmentations across 1.0K sequences accompanied by 37 object models. To augment spacial relation between human and object, we present an innovative collaboration retargeting technique in Section 3.3. This technique integrates CORE4D-Real with CORE4D-Synthetic, thereby expanding our collection with an additional 10K sequences and 3K rigid objects. Detailed characteristics such as data diversities are discussed in Section 3.4.
3.1 CORE4D-Real Data Capture
Figure 1: CORE4D-Real data capturing system. (a) demonstrates the wearing of mocap suits and the positioning of the egocentric camera. (b) shows an object with four markers. (c) illustrates the data capturing system and camera views.
Figure 2: CORE4D-Real data modality overview.
To collect precise human-object motions with visual signals, we set up a hybrid data capturing system shown in Fig. 1, consisting of an inertial-optical mocap system, four allocentric RGB-D cameras and a camera worn by participants for egocentric sensing. The system operates at 15 FPS.
Inertial-optical Mocap System. To accurately capture human-object poses in multi-person collaboration scenarios, often involving severe occlusion, we use an inertial-optical mocap system [74] inspired by CHAIRS [41] This system includes 12 infrared cameras, mocap suits with 8 inertial-optical trackers and two data gloves per person, and markers of a 10mm radius. The mocap suits capture Biovision Hierarchy (BVH) skeletons of humans, while markers attached to the objects track object motion.
Visual Sensors. Kinect Azure DK cameras are integrated to capture allocentric RGB-D signals, and an Osmo Action3 is utilized to capture egocentric color videos. The resolution of all the visual signals is 1920x1080. Cameras are calibrated by the mocap system and synchronized via timestamp. Details on camera calibration and synchronization are in the appendix.
Object Model Acquisition. CORE4D-Real includes 37 3D models of rigid objects spanning six household object categories. Each object model is constructed by an industrial 3D scanner with up to 100K triangular faces. We additionally adopt manual refinements on captured object models to remove triangle outliers and improve accuracy.
Privacy Protection. To ensure participant anonymity, blurring is applied to faces [76] in RGB videos. The participants all consented to releasing CORE4D, and were notified of their right to remove their data from CORE4D at any time.
3.2 CORE4D-Real Data Annotation
Object Pose Tracking. To acquire the 6D pose of a rigid object, we attach four to five markers to the object’s surface. The markers formulate a virtual rigid that the mocap system can track. With accurate localization of the object manually, the object pose can be precisely determined by marker positions captured by the infrared cameras.
Human Mesh Acquisition. Aligning with existing datasets [41, 53], we retarget BVH [67] human skeletons to SMPL-X [79]. SMPL-X [79] formulates a human mesh as $D_{\text{smplx}}=M(\beta,\theta)$. The body shape $\beta\in\mathbb{R}^{10}$ are optimized to fit the constraints on manually measured human skeleton lengths. With $\beta$ computed, we optimize the full-body pose $\theta\in\mathbb{R}^{159}$ with the loss function:
$\mathcal{L}=\mathcal{L}_{\text{reg}}+\mathcal{L}_{j\text{3D}}+\mathcal{L}_{j\text{Ori}}+\mathcal{L}_{\text{smooth}}+\mathcal{L}_{h\text{3D}}+\mathcal{L}_{h\text{Ori}}+\mathcal{L}_{\text{contact}},$ (1)
where $\mathcal{L}_{\text{reg}}$ ensures the simplicity of the results and prevents unnatural, significant twisting of the joints. $\mathcal{L}_{j\text{3D}}$ and $\mathcal{L}_{j\text{Ori}}$ encourage the rotation of joints and the global 3D positions to closely match the ground truth. $\mathcal{L}_{h\text{3D}}$ and $\mathcal{L}_{h\text{Ori}}$ guide the positioning and orientation of the fingers. $\mathcal{L}_{\text{smooth}}$ promotes temporal smoothness. $\mathcal{L}_{\text{contact}}$ encourages realistic contact between the hands and objects. Then using SMPL-X [79] $M(\beta,\theta,\Phi):\mathbb{R}^{|\theta|\times|\beta|}\mapsto\mathbb{R}^{3N}$ to generate human mesh. Details on loss functions are in the appendix.
2D Mask Annotation. We offer automatic 2D segmentation for individuals and the manipulated objects to aid in predictive tasks like vision-based human-object pose estimation [3, 114]. We first use DEVA [10] to segment human and object instances in a captured interaction image with text prompts. Then, we render human and object meshes separately on each image and select the instance with the highest Intersection-over-Union (IoU) for mask annotation.
Figure 3: Collaboration retargeting pipeline. We propose a collaboration retargeting algorithm by iteratively refining interaction motion. The input is a source-target pair. First, we sample contact candidates from whole CORE4D-Real contact knowledge on source. For each contact candidate, we apply contact retargeting to propagate contact candidates to contact constraints on target. Sampled motion from CORE4D-Real provides a high-level collaboration pattern, together with low-level contact constraints, we obtain interaction candidates from interaction retargeting. Then, the human pose discriminator selects the optimal candidates, prompting a contact constraints update via beam search. After multiple iterations, the process yields augmented interactions. This iterative mechanism can effectively get a reasonable one from numerous contact constraints and ensures a refined interaction, enhancing the dataset’s applicability across various scenarios.
3.3 CORE4D-Synthetic Data Generation
In order to enrich the diversities of object geometries and human-object spatial relations, our retargeting algorithm transfers real interactions to ShapeNet [7] objects of the same category, thereby significantly expanding the dataset regarding the object’s diversity. When transferring interactions across objects, contact points are always the key and it is important to consider whether they can be properly transferred with consistent semantics on new objects [127, 144]. However, we find this insufficient when object geometries vary largely and correspondences become hard to build. We thus tackle interaction retargeting from a novel human-centric perspective where good contact points should support natural human poses and motions. We realize this idea through the pipeline depicted in Figure 3, which comprises three key components. First, object-centric contact retargeting uses whole contact knowledge from CORE4D-Real to obtain accurate contact with different objects. Second, contact-guided interaction retargeting adapts motion sequences to new object geometries while considering the contact constraints. Third, a human-centric contact selection evaluates poses from interaction candidates to select the most plausible contacts.
Object-centric Contact Retargeting. To acquire reasonable human poses, contact constraints on the target object are essential. We draw inspiration from Tink [128] and train DeepSDF on all objects’ signed distance fields (SDFs). For source object SDF $O_{s}$ and target object SDF $O_{t}$, we first apply linear interpolation on their latent vectors $o_{s}$ and $o_{t}$ and obtain $N$ intermediate vectors $o_{i}=\frac{N+1-i}{N+1}o_{s}+\frac{i}{N+1}o_{t}(1\leq i\leq N)$. We then decode $o_{i}$ to its SDF $O_{i}$ via the decoder of DeepSDF, and reconstruct the corresponding 3D mesh $M_{i}$ using the Marching Cubes algorithm [61]. Thereby get mesh sequence $\mathcal{M}=[\textit{source},M_{1},M_{2},...,M_{N},\textit{target}]$ and successively transfer contact positions between every two adjacent meshes in $\mathcal{M}$ via Nearest-neighbor searching. In addition, we leverage all contact candidates from CORE4D-Real on source to form a pool of contact candidates and transfer them to target as contact constraints.
Contact-guided Interaction Retargeting. For each contact constraint, interaction retargeting aims to transfer human interaction from source to target. To greatly enforce the consistency of interaction motion, we optimize variables including the object rotations $R_{o}\in\mathbb{R}^{N\times 3}$ and translations $T_{o}\in\mathbb{R}^{N\times 3}$, human poses $\theta_{1,2}\in\mathbb{R}^{2\times N\times 153}$, translation $T_{1,2}\in\mathbb{R}^{2\times N\times 3}$ and orientation $O_{1,2}\in\mathbb{R}^{2\times N\times 3}$ on the SMPL-X [79]. $N$ is the frame number.
We first estimate the target’s motion $\{R_{o},T_{o}\}$ by solving an optimization problem as follows:
$R_{o},T_{o}\longleftarrow\mathop{\operatorname{argmin}}\limits_{R_{o},T_{o}}(\mathcal{L}_{f}+\mathcal{L}_{\text{spat}}+\mathcal{L}_{\text{smooth}}),$ (2)
where fidelity loss $\mathcal{L}_{f}$ evaluates the difference of the target’s rotation and translation against the source, restriction loss $\mathcal{L}_{\text{spat}}$ penalizes target’s penetration with the ground, and smoothness loss $\mathcal{L}_{\text{smooth}}$ constrains the target’s velocities between consecutive frames.
Given the target’s motion and contact constraints, we then transfer humans’ interactive motion $\{\theta_{1,2},T_{1,2},O_{1,2}\}$ from the source to the target by solving another optimization problem as follows:
$\theta_{1,2},T_{1,2},O_{1,2}\longleftarrow\mathop{\operatorname{argmin}}\limits_{\theta_{1,2},T_{1,2},O_{1,2}}(\mathcal{L}_{j}+\mathcal{L}_{c}+\mathcal{L}_{\text{spat}}+\mathcal{L}_{\text{smooth}}),$ (3)
where fidelity loss $\mathcal{L}_{\text{j}}$ evaluates the difference in human joint positions before and after the transfer, contact loss $\mathcal{L}_{c}$ computes the difference between human-object contact regions and the contact constraints, $\mathcal{L}_{\text{spat}}$ and $\mathcal{L}_{\text{smooth}}$ ensures the smoothness of human motion. Details on the loss designs are in the appendix.
Human-centric Contact Selection. Selecting reasonable contact constraints efficiently is challenging due to their large scales and the time-consuming interaction retargeting. We address this challenge by developing a beam search algorithm to select contact constraints from a human-centric perspective. Specifically, we train a human pose discriminator inspired by GAN-based motion generation works [125, 118]. To train it, we build a pairwise training dataset, with each pair consisting of one positive human pose sample and one negative one. Positive samples are encouraged to get higher scores than negative ones. We use CORE4D-Real as positive samples. We add 6D pose noise $\Delta(\alpha,\beta,\gamma,x,y,z)$ on target motion, and regard corresponding human motions generated by contact-guided interaction retargeting as negative samples. The loss function is:
$\mathcal{L}_{\text{ranking}}=-\log(\sigma(R_{\text{pos}}-R_{\text{neg}}-m(S_{\text{pos}},S_{\text{neg}}))),$ (4)
where $S_{\text{pos}}$ and $S_{\text{neg}}$ denote inputs for positive and negative samples respectively, with $R_{\text{pos}}$ and $R_{\text{neg}}$ being their corresponding discriminator scores. $\sigma$ is Sigmoid function, and $m(S_{\text{pos}},S_{\text{neg}})=||\Delta(\alpha,\beta,\gamma,x,y,z)||$ is human-guide margin [77] between positive and negative poses. This margin could explicitly instruct the discriminator to yield more significant disparities across different poses.
To ensure the realism of human interactions, we also introduce an interpenetration penalty. We prioritize those with the highest discriminator scores while ensuring acceptable levels of interpenetration as the optimal contact constraints.
3.4 Dataset Characteristics
Figure 4: Dataset statistics. (a) shows object samples from six categories. Bars in (b) indicate when the person is in contact with the object during the entire collaborative object rearrangement interaction process. (c) presents the proportion of collaboration modes in the dataset.
To better model collaborative object rearrangement interactions, we focus on diversifying our dataset in several vital areas: object geometries, collaboration modes, and 3D scenes. These ensure a comprehensive representation of real-world interactions.
Diversity in Object Geometries. We design six object categories to cover the main collaborative object rearrangement interaction scenarios as Fig. 4(a). Categories with relatively simple geometry, uniformity, and typically exhibiting symmetry include box, board, barrel, and stick. Categories with more complex geometries and significant individual differences include chair and desk.
Diversity in Collaboration Modes. We define five human-human collaboration modes in collaborative object rearrangement. Each mode represents a unique form of collaboration between two individuals, providing a new perspective and possibilities for understanding and researching collaborative behaviors. At first, we define the person with the egocentric camera as Person 2, and the other as Person 1. Collaborative carrying tasks are divided by whether Person 2 knows the goal or not. Tasks of handover and solely move alternate between the two participants. In join and leave tasks, Person 2 will either join in to help or leave halfway through, respectively.
Diversity in 3D Scenes. Surrounding scenarios are set up with varying levels of scene complexity: no obstacle, single obstacle, and many obstacles (more than one). Participants are asked to navigate through these randomly placed obstacles by their own means. We observe that this typically involved behaviors including bypassing, going through, stepping over, or moving obstacles aside.
4 Experiments
In this section, we first present the train-test split of CORE4D (Section 4.1). We then propose two benchmarks for generating human-object collaboration: human-object motion forecasting (Section 4.2), and interaction synthesis (Section 4.3). Finally, Section 4.4 presents extensive studies on the collaboration retargeting approach.
4.1 Data Split
We construct a training set from a random assortment of real objects, combining their real motions and corresponding synthetic data. We also create two test sets from CORE4D-Real for non-generalization and inner-category generalization studies. Test set S1 includes interactions with training set objects, while S2 features interactions with new objects. CORE4D-Synthetic is not included in the test set, avoiding potential biases from the retargeting algorithm. Details are shown in the appendix.
4.2 Human-object Motion Forecasting
Forecasting 4D human motion [28, 80, 65, 29] is a crucial problem with applications in VR/AR and embodied perception [44]. Current research [15, 1, 103, 119] is limited to individual behaviors due to data constraints. Our work expands this by using diverse multi-person collaborations, making the prediction problem both intriguing and challenging.
Task Formulation. Given the object’s 3D model and human-object poses in adjacent 15 frames, the task is to predict their subsequent poses in the following 15 frames. The human pose $P_{h}\in\mathbb{R}^{23\times 3}$ represents joint rotations of the SMPL-X [79] model, while the object pose $P_{o}=\{R_{o}\in\mathbb{R}^{3},T_{o}\in\mathbb{R}^{3}\}$ denotes 3D orientation and 3D translation of the rigid object model.
Evaluation Metrics. Following existing motion forecasting works [14, 102, 119], we evaluate human joints position error $J_{e}$, object translation error $T_{e}$, object rotation error $R_{e}$, human-object contact accuracy $C_{\text{acc}}$, and penetration rate $P_{r}$. Details are provided in the appendix.
Methods, Results, and Analysis. We evaluate three state-of-the-art motion forecasting methods, MDM [97], InterDiff [119], and CAHMP [14]. Table 2 quantitatively shows these methods reveal a consistent drop in performance for unseen objects (S2) versus seen ones (S1) regarding human pose prediction. Meanwhile, errors in object pose prediction remain similar. This highlights the challenges in generalizing human collaborative motion for novel object shapes.
Test Method Human Object Contact
Set $J_{e}$ ($\downarrow$) $T_{e}$ ($\downarrow$) $R_{e}$ ($\downarrow$) $C_{\text{acc}}$ ($\uparrow$) $P_{r}$ ($\downarrow$)
S1 MDM [97] 170.8 136.8 10.7 84.9 0.3
InterDiff [119] 170.8 135.1 10.2 84.9 0.3
CAHMP [14] 169.4 110.3 9.0 - -
S2 MDM [97] 186.4 136.0 11.1 88.0 0.3
InterDiff [119] 186.4 133.6 10.7 88.0 0.3
CAHMP [14] 170.5 112.9 9.5 - -
Table 2: Quantitative results on motion forecasting.
4.3 Interaction Synthesis
Generating human-object interaction [53, 22, 52, 81] is an emerging research topic benefiting human avatar animation and human-robot collaboration [12, 73]. With extensive collaboration modes and various object categories, CORE4D constitutes a knowledge base for studying generalizable algorithms of human-object-human interactive motion synthesis.
Task Formulation. Following recent studies [93, 53], we define the task as object-conditioned human motion generation. Given an object geometry sequence $G_{o}\in\mathbb{R}^{T\times N\times 3}$, the aim is to generate corresponding two-person collaboration motions $M_{h}\in\mathbb{R}^{2\times T\times 23\times 3}$. This involves frame numbers $T$, object point clouds $G_{o}$, and human pose parameters for the SMPL-X [79] model.
Evaluation Metrics. Following individual human-object interaction synthesis [53], we evaluate human joint position error $R.J_{e}$, object vertex position error $R.V_{e}$, and human-object contact accuracy $C_{\text{acc}}$. The FID score (FID) is leveraged to quantitatively assess the naturalness of synthesized results. Details of the metric designs are presented in the appendix.
Methods, Results, and Analysis. We utilize three advanced diffusion models [97, 53, 52] as baselines. MDM [97] and CHOIS [52] are one-stage conditional motion diffusion models, while OMOMO is a two-stage approach with hand positions as intermediate results. Quantitative evaluations reveal larger errors in OMOMO when modeling multi-human collaboration compared to individual interaction synthesis by Li et al. [53]. Furthermore, the synthesized results have a higher FID than real motion data, indicating challenges in motion naturalness.
Test Set Method $R.J_{e}$ ($\downarrow$) $R.V_{e}$ ($\downarrow$) $C_{\text{acc}}$ ($\uparrow$) FID ($\downarrow$)
S1 MDM [97] 138.3 194.8 76.5 7.5
OMOMO [53] 138.0 196.9 78.0 7.8
CHOIS [52] 138.4 194.3 76.2 7.7
S2 MDM [97] 146.1 208.3 76.6 7.9
OMOMO [53] 145.3 209.9 77.8 7.4
CHOIS [52] 145.8 206.7 76.2 7.7
Table 3: Quantitative results on interaction synthesis.
4.4 Collaboration Retargeting
Comparisons Designs Phys. Eval. User Preferences
C D U $P$ $C_{\text{acc}}$ Contact Motion
($\downarrow$) ($\uparrow$) A / B / Approx. ($\uparrow$)
Abl.1 A 0.61 83.2 7.8/88.9/3.3 3.3/84.4/12.3
B 0.24 83.3
Abl.2 A 0.55 82.9 1.2/91.4/7.4 3.2/85.1/11.7
B 0.24 83.3
Abl.3 A * 0.68 94.7 5.6/84.3/10.1 2.2/86.6/11.2
B 0.24 83.3
Abl.4 A 0.24 83.3 5.0/76.0/19.0 4.0/69.0/27.0
B 0.23 85.5
Table 4: Ablation study. C, D, and U denote contact candidates, the human pose discriminator, and the contact candidate update, respectively. $P$ is penetration distance. $C_{acc}$ is contact accuracy.
User Studies. We conduct user studies to examine the quality of CORE4D-Synthetic in terms of naturalness of contact and human motion. Each study comprises two collections, each with at least 100 sequences displayed in pairs on a website. Users are instructed to assess the realism of human-object contacts and the naturalness of human motions, and then select the superior one in each pair separately. Recognizing the diversity of acceptable contacts and motions, participants are permitted to deem the performances as roughly equivalent.
Ablation on Contact Candidates. In Table 4.Abl.1, we only use the contact points from a source trajectory for retargeting to the target instead of resorting to the CORE4D-Real for many candidates, making the whole retargeting process similar to the OakInk [128] method. We observe a sharp decline in both physical plausibility and user preferences, indicating that our method compensates for OakInk’s shortcomings in retargeting objects with significant geometric and scale variations.
Ablations on Discriminator. In Table 4.Abl.2, we omit the human pose discriminator in the collaboration retargeting. Method A randomly chooses a candidate from the contact candidates. There are obvious performance drops, demonstrating the critical role of the human pose discriminator in selecting appropriate candidates. Table 4.Abl.3 further compares the proposed discriminator against selecting the motion with the most accurate contact (method A), and user studies reveal significant superiority of the discriminator.
Ablation on Contact Candidate Update. We exclude contact candidate update process in Table 4.Abl.4 experiment. This removal has weakened our method’s ability to search for optimal solutions on objects, resulting in a modest degradation in penetration distance. The user study still exhibited a strong bias, indicating a perceived decline in the plausibility of both contact and motion. This ablation underscores the importance of contact candidate updates within our methodology.
Comparing CORE4D-Synthetic with CORE4D-Real. We assess the quality of CORE4D-Synthetic by comparing it with CORE4D-Real through user study. In conclusion, there is a 43% probability that users perceive the quality of both options as comparable. Furthermore, in 14% of cases, users even exhibit a preference for synthetic data. This indicates that the quality of our synthetic data closely approximates that of real data.
5 Dataset Applications
5.1 CORE4D-Synthetic Enhances Human-object Motion Forecasting Quality
Table 5 compares the motion forecasting ability of light-weighted CAHMP [16]. The test set is S2 defined in Section 4.1. We assess the quality of CORE4D-Synthetic by comparing No.A and No.B. No.A even have better performance on object due to enriched spacial relation between human and object in CORE4D-Synthetic. No.C shows the value of the CORE4D-Synthetic by largely improving the performance. Details are in the appendix.
No Train Set Human Object
Total Real Synthetic $J_{e}$ ($\downarrow$) $T_{e}$ ($\downarrow$) $R_{e}$ ($\downarrow$)
A 1.0K 0.1K 0.9K 127.7 121.7 8.04
B 1.0K 1.0K 0 127.0 120.5 9.48
C 5.0K 1.0K 4.0K 116.2 112.1 6.99
Table 5: Ablation on the incorporation of CORE4D-Synthetic on the motion forecasting task.
5.2 CORE4D Supports Humanoid Skill Learning
Benefitting from rapid developments of humanoid robots [25, 47, 100], tremendous progress has been made in studying versatile humanoid skills for locomotion [35, 94, 84, 66] and humanoid-object interaction [68, 21, 133]. Aiming at enabling humanoid learning skills from human data, we select human interaction motions with three large boxes from CORE4D, and retarget them onto the Unitree H1 humanoid robot [100] with object scale augmentation. With 890 human-like humanoid-box interaction sequences, we design a box-lifting task in Isaac Gym [64], and benchmark two state-of-the-art humanoid imitation learning (IL) methodologies [143, 26] comparing to demonstration-free reinforcement learning (RL) paradigm [87].
Table 6 compares the success rates of these methods. Leveraging interaction data from CORE4D, the two IL methods [143, 26] consistently make it possible for humanoids to lift unseen boxes with visual sensor signals successfully, demonstrating that CORE4D can promote humanoid interaction skill learning. Figure 5 exemplifies a successful case of HumanPlus [26], showing that humanoids can learn from CORE4D and achieve the task in a human-like manner. As the development of multi-humanoid imitation learning methods in the future, we anticipate that CORE4D can further promote collaboration skill learning. Details on task formulation, method designs, and evaluations are in the appendix.
Data N/A CORE4D
Method PPO [87] HumanPlus [26] HST [26] + ACT [143]
$SR$ ($\uparrow$) 0.0 21.0 26.5
Table 6: Success rates of RL and IL in humanoid box lifting.
Figure 5: Visualization of the humanoid box-lifting skill trained by CORE4D via imitation learning.
6 Conclusion and Limitations
We present CORE4D, a novel large-scale 4D human-object-human interaction dataset for collaborative object rearrangement. It comprises diverse compositions of various object geometries, collaboration modes, and surrounding 3D scenes. To efficiently enlarge the data scale, we contribute a hybrid data acquisition method involving real-world data capturing and a novel synthetic data augmentation algorithm, resulting in 11K motion sequences covering 37 real-world and 3K virtual objects. Extensive experiments demonstrate the effectiveness of the data augmentation strategy and the value of the augmented motion data. We benchmark human-object motion forecasting and interaction synthesis on CORE4D, revealing new challenges and research opportunities.
Limitations. Firstly, outdoor scenes are not incorporated in CORE4D-Real due to the usage of the mocap system. Secondly, visual signals are excluded in CORE4D-Synthetic. Transferring real-world videos onto synthesized collaboration motions could be an interesting future direction.
Appendix
Contents:
  • A. Cross-dataset Evaluation
  • B. Details on Real-world Data Acquisition
  • C. Details on CORE4D-Synthetic Data Generation
  • D. Dataset Statistics and Visualization
  • E. Details on Data Split
  • F. Evaluation Metrics for Benchmarks
  • G. Qualitative Results on Benchmarks
  • H. Details on the Application of CORE4D-Synthetic
  • I. Details on Humanoid Skill Learning using CORE4D
  • J. CORE4D-Real Data Capturing Instructions
Appendix A Cross-dataset Evaluation
To examine the data quality of CORE4D-Real, we follow existing dataset efforts [123, 8, 60] and conduct the vision-based cross-dataset evaluation. We select an individual human-object-interaction dataset BEHAVE [3] that includes color images and select 2D human keypoint estimation as the evaluation task.
Data Preparation. For a color image from CORE4D-Real and BEHAVE [3], we first detect the bounding box for each person via ground truth human pose and obtain the image patch for the person. We then resize the image patch to get a maximal length of 256 pixels and fill it up into a 256x256 image with the black color as the background. Finally, for each 256x256 image, we automatically acquire the ground truth 2D-pixel coordinates of 22 SMPL-X [79] human body joints from 3D human poses. For data split, we follow the original train-test split for BEHAVE [3] and merge the two test sets (S1, S2) for CORE4D-Real.
Task Formulation. Given a 256x256 color image including a person, the task is to estimate the 2D-pixel coordinate for each of the 22 SMPL-X [79] human body joints.
Evaluation Metrics. $P_{e}$ denotes the mean-square error of 2D coordinate estimates. $Acc$ denotes the percentage of the coordinate estimates with the Euclidean distance to the ground truth smaller than $15$ pixels.
Method, Results, and Analysis. We draw inspiration from HybrIK-X [51] and adopt their vision backbone as the solution. Table 7 shows the method performances on the two datasets under different training settings. Due to the significant domain gaps in visual patterns and human behaviors, transferring models trained on one dataset to the other would consistently encounter error increases. Despite the domain gaps, integrally training on both datasets achieves large performance gains on both CORE4D-Real and BEHAVE [3], indicating the accuracy of CORE4D-Real and the value of the dataset serving for visual perception studies.
Test Train CORE4D-Real BEHAVE [3] CORE4D-Real +BEHAVE [3]
CORE4D-Real 152.4 / 91.2 904.9 / 35.6 121.7 / 92.4
BEHAVE[3] 887.9 / 37.8 146.3 / 88.9 128.2 / 89.8
[…]
Appendix B Details on Real-world Data Aquisition
In this section, we describe our system calibration (Section B.1) and time synchronization (Section B.2) in detail. Moreover, we provide detailed information on loss functions of the human mesh acquisition (Section B.3).
B.1 System Calibration
Calibrating the Inertial-optical Mocap System. Three reflective markers are fixed at known positions on a calibration rod, by which the 12 high-speed motion capture cameras calculate their relative extrinsic parameters, providing information about their spatial relationships. Additionally, three markers fixed at the world coordinate origin are employed to calibrate the motion capture system coordinate with the defined world coordinate.
Calibrating Camera Intrinsic. The intrinsic parameters of allocentric and egocentric cameras are calibrated using a chessboard pattern.
Calibrating Extrinsic of the Allocentric Cameras. We place ten markers in the camera view to locate each allocentric camera. By annotating the markers’ 3D positions in the world coordinate system and their 2D-pixel coordinates on allocentric images, the camera’s extrinsic parameters are estimated by solving a Perspective-n-Point (PnP) problem via OpenCV.
Calibrating Extrinsic of the Egocentric Camera. We obtain the camera’s pose information by fixing the camera to the head tracker of the motion capture suit. Similarly, ten markers are used to calibrate the relative extrinsic parameters of the first-person perspective cameras, allowing for determining their positions and orientations relative to the motion capture system. Additionally, to mitigate errors introduced by the integration of optical and inertial tracking systems, a purely optical tracking rigid is mounted on the motion camera.
B.2 Time Synchronization
To implement our synchronization method, we first set up a Network Time Protocol (NTP) server on the motion capture host. This server serves as the time synchronization reference for the Windows computer connected to the Kinect Azure DK. We minimize time discrepancies by connecting the Windows computer to the NTP server in high-precision mode and thus achieving precise synchronization.
Additionally, we employ a Linear Timecode (LTC) generator to encode a time signal onto the action camera’s audio track. This time signal serves as a synchronization reference for aligning the first-person perspective RGB information with the motion capture data.
B.3 Loss Function Designs for Human Mesh Acquisition
To transfer the BVH [67] human skeleton to the widely-used SMPL-X [79] model. We optimize body shape parameters $\beta\in\mathbb{R}^{10}$ to fit the constraints on manually measured human skeleton lengths and then optimize the full-body pose $\theta\in\mathbb{R}^{159}$ with the following loss function:
$\mathcal{L}=\mathcal{L}_{\text{reg}}+\mathcal{L}_{j\text{3D}}+\mathcal{L}_{j\text{Ori}}+\mathcal{L}_{\text{smooth}}+\mathcal{L}_{h\text{3D}}+\mathcal{L}_{h\text{Ori}}+\mathcal{L}_{\text{contact}}.$ (5)
Regularization Loss $\mathcal{L}_{\text{reg}}$. The regularization loss term is defined as
$\mathcal{L}_{\text{reg}}=\sum\left|\left|\theta_{\text{body}}\right|\right|^{2}\cdot\lambda_{\text{body}}+\left(\sum\left|\left|\theta_{l\_\text{hand}}\right|\right|^{2}+\sum\left|\left|\theta_{r\_\text{hand}}\right|\right|^{2}\right)\cdot\lambda_{\text{hand}},$ (6)
where $\theta_{\text{body}}\in\mathbb{R}^{21\times 3}$ represents the body pose parameters defined by 21 joints of the skeleton, $\theta_{l\_hand}\in\mathbb{R}^{12}$ and $\theta_{r\_\text{hand}}\in\mathbb{R}^{12}$ represents the hand pose parameters. For each hand, the original SMPL-X skeleton has 15 joints with parameters $\theta_{\text{hand}}\in\mathbb{R}^{15\times 3}$. However, principal component analysis (PCA) is applied to the hand pose parameters. The $\theta_{\text{hand}}$ parameters are transformed into a lower-dimensional space, specifically $\mathbb{R}^{12}$. $\lambda_{\text{body}}=10^{-3}$ and $\lambda_{\text{hand}}=10^{-4}$ are different weights that are used to control the regularization strength for the body and hand pose parameters, respectively. This loss ensures the simplicity of the results and prevents unnatural, significant twisting of the joints.
3D Position Loss $\mathcal{L}_{j\text{3D}}$ and $\mathcal{L}_{h\text{3D}}$. The 3D position loss term is defined as
$\mathcal{L}_{\text{3D}}=\sum\left|\left|\textbf{T}_{\text{smplx}}-\textbf{T}_{\text{bvh}}\right|\right|^{2}\cdot\lambda_{\text{3D}},$ (7)
where $\textbf{T}_{\text{smplx}}\in\mathbb{R}^{3}$ represents the 3D global coordinates of the joints in the SMPL-X model and $\textbf{T}_{\text{bvh}}\in\mathbb{R}^{3}$ represents the corresponding 3D global coordinates of the joints in the BVH representation. $\mathcal{L}_{j\text{3D}}$ represents the 3D position loss sum for the 21 body joints, while $\mathcal{L}_{h\text{3D}}$ represents the 3D position loss sum for the 30 hand joints (15 joints per hand). These two terms have different weights, set as $\lambda_{j\text{3D}}=1.0$ and $\lambda_{h\text{3D}}=2.0$, respectively.
Orientation Loss $\mathcal{L}_{j\text{Ori}}$ and $\mathcal{L}_{h\text{Ori}}$. The orientation loss term is defined as
$\mathcal{L}_{\text{Ori}}=\sum\left|\left|\textbf{R}_{\text{smplx}}-\textbf{R}_{\text{bvh}}\right|\right|^{2}\cdot\lambda_{\text{Ori}},$ (8)
which is similar to $\mathcal{L}_{\text{3D}}$, except that $\mathcal{R}_{\text{smplx}}\in\mathbb{R}^{3\times 3}$ and $\mathcal{R}_{\text{bvh}}\in\mathbb{R}^{3\times 3}$ represent the rotation matrices for the adjacent joints in the SMPL-X and corresponding BVH representations, respectively. Specifically, body joints named head, spine, spine2, leftUpLeg, rightUpLeg, rightShoulder, leftShoulder, rightArm, leftArm, and neck are subjected to orientation loss, ensuring that their rotations relative to adjacent nodes are close to the BVH ground truth. $\lambda_{\text{Ori}}$ is set to $0.2$.
Temporal Smoothness Loss $\mathcal{L}_{\text{smooth}}$. The temporal smoothness loss term is defined as
$\mathcal{L}_{\text{smooth}}=\sum_{i=1}^{N}\left(\left|\left|\theta_{i}-\theta_{i-1}\right|\right|^{2}\right)\cdot\lambda_{\text{smooth}}$ (9)
where $\theta_{i}\in\mathbb{R}^{(21+30)\times 3}$ represents the body and hand pose of the $i$-th frame. $\lambda_{\text{smooth}}$ is set to $20.0$.
Contact Loss $\mathcal{L}_{\text{contact}}$. The contact loss term is defined as
$\mathcal{L}_{\text{contact}}=\sum\left(\left|\left|\textbf{T}_{\text{finger}}-\textbf{T}_{\text{obj}}\right|\right|^{2}\cdot\mathcal{J}(\textbf{T}_{\text{finger}},\textbf{T}_{\text{obj}})\right)\cdot\lambda_{\text{contact}}$ (10)
where $\mathcal{T}_{\text{finger}}\in\mathbb{R}^{10\times 3}$ is the global coordinates of ten fingers, and $\mathcal{T}_{\text{obj}}\in\mathbb{R}^{10\times 3}$ is the corresponding global coordinates of the point closest to finger. $\mathcal{J}(\textbf{T}_{\text{finger}},\textbf{T}_{\text{obj}})$ is 1 when the distance between $\textbf{T}_{\text{finger}}$ and $\textbf{T}_{\text{obj}}$ is less than a threshold, otherwise it is 0. And $\lambda_{\text{contact}}$ is $2.0$.
Appendix C Details on CORE4D-Synthetic Data Generation
In this section, we provide details on our synthetic data generation (collaboration retargeting) method. Firstly, we clarify term definitions in Section C.1. We then explicitly introduce the whole method pipeline in detail in Section C.2. Finally, we provide implementation details in Sections C.3 and C.4.
C.1 Term Definitions
We provide definitions for the terms in our collaboration retargeting pipeline as follows.
Contact Candidate: Contact candidate is a quadruple list containing all possible contact region index (person1_leftHand, person1_rightHand, person2_leftHand, person2_rightHand) on source’s vertices. For each source, we record the contact regions of the four hands in each frame of each data sequence. At the beginning of the synthetic data generation pipeline, we sample contact candidates from these records.
Contact Constraint: Having contact candidate on source, we apply DeepSDF-based [78] contact retargeting to transfer the contact regions to target. These contact regions on target are the contact constraints fed into the contact-guided interaction retargeting module.
Source Interaction: During each collaboration retargeting process, we sample a human-object-human collaborative motion sequence from CORE4D-Real as the source interaction to guide temporal collaboration pattern.
Interaction Candidate: Sampling $N$ contact candidates, we apply contact-guided interaction retargeting $N$ times and have $N$ human-object-human motion outputs, dubbed interaction candidates. These motions would be fed into the human-centric contact selection module to assess their naturalness.
C.2 Method Pipeline
The algorithm takes a source-target pair as input. First, we sample contact candidates from the whole CORE4D-Real contact knowledge on source. For each contact candidate, we apply object-centric contact retargeting to propagate contact candidates to contact constraints on target. Sampling motion from CORE4D-Real provides a high-level temporal collaboration pattern, and together with augmented low-level spatial relations, we obtain interaction candidates from the contact-guided interaction retargeting. Then, the human-centric contact selection module selects the optimal candidates, prompting a contact constraint update. After multiple iterations, the process yields augmented interactions. This iterative mechanism ensures a refined augmentation of interactions, enhancing the dataset’s applicability across various scenarios.
C.3 Contact-guided Interaction Retargeting
The contact-guided interaction retargeting is a two-step optimization. We start by optimizing the motion of target. Then with target contact constraints, we optimize the poses of the two persons.
Object motion retargeting. We deliberately design temporal and spatial losses to acquire consistent and smooth target motion. In the concern of efficiency, we jointly optimize all frames in a single data sequence with $N$ frames. To guarantee the fidelity of object motion, we design the fidelity loss $L_{f}$ to restrict the rotation $R_{o,i}$ and the translation $T_{o,i}$ with the ground-truth rotation $R^{\prime}_{o,i}$ and translation $T^{\prime}_{o,i}$ in $N$ frames:
$\mathcal{L}_{f}=\lambda_{f}\sum\limits_{i}(||R^{\prime}_{o,i}-R_{o,i}||_{1}+||T^{\prime}_{o,i}-T_{o,i}||_{1}).$ (11)
We then address restriction on target’s spatial position to avoid penetration with the ground. The spatial loss is defined as:
$\mathcal{L}_{\text{spat}}=\lambda_{\text{spat}}\sum\limits_{i}\text{max}(-\text{min}(\text{height}_{i}),0),$ (12)
where $\text{min}(\text{height}_{i})$ represents the lowest spatial position of the objects per frame. A smoothness loss is designed to constrain the object pose difference between consecutive frames:
$\mathcal{L}_{\text{smooth}}=\lambda_{\text{smooth}}\sum\limits_{i}a_{R_{o,i}}^{2}+a_{T_{o,i}}^{2},$ (13)
where $a$ is the acceleration of rotation and translation during $N$ frames defined as:
$a_{R_{o,i}}=2R_{o,i}-R_{o,i-1}-R_{o,i+1},$ (14)
$a_{T_{o,i}}=2T_{o,i}-T_{o,i-1}-T_{o,i+1},$ (15)
The total object motion retargeting problem is:
$R_{o},T_{o}\longleftarrow{}\mathop{\operatorname{argmin}}\limits_{R_{o},T_{o}}(\mathcal{L}_{f}+\mathcal{L}_{\text{spat}}+\mathcal{L}_{\text{smooth}}).$ (16)
Human motion retargeting. We next optimize each person’s motion based on the motion of target and the contact constraint. To acquire visually plausible motion, we design the fidelity loss $\mathcal{L}_{j}$ and the smoothness loss $\mathcal{L}_{\text{smooth}}$. Besides, we utilize the contact correctness loss $\mathcal{L}_{c}$ to acquire contact consistency in target interaction motion, and leverage spatial loss $L_{\text{spat}}$ similar to Equation 12 to avoid human-ground inter-penetration.
To enhance motion fidelity, we define two loss functions $\mathcal{L}_{\text{sr}}$ and $\mathcal{L}_{\text{wr}}$ and let $L_{j}=\mathcal{L}_{\text{sr}}+\mathcal{L}_{\text{wr}}$. For joints from the human arms, despite following the correct temporal collaboration pattern, their global positions would vary concerning diverse object geometries. Therefore, we utilize oriented vectors pointing to their parent body joints to obtain a relative joint fidelity:
$\mathcal{L}_{\text{sr}}=\lambda_{\text{sr}}\sum\limits_{i}\sum\limits_{j\in\text{arm}}\|(P_{j,i}-P_{\text{parent}(j),i})-(P^{\prime}_{j,i}-P^{\prime}_{\text{parent}(j),i})\|_{2}^{2},$ (17)
where $P_{j,i}$ denotes the 3D global position of joint $j$ in frame $i$, and $P^{\prime}$ denotes ground-truth values. $\mathcal{L}_{\text{wr}}$ denotes constraints on the global positions of other joints:
$\mathcal{L}_{\text{wr}}=\lambda_{\text{wr}}\sum\limits_{i}\sum\limits_{j\notin\text{arm}}\|P_{j,i}-P^{\prime}_{j,i}\|_{2}^{2}.$ (18)
The design of the smoothness loss is similar to Equation 13, penalizing huge acceleration of human SMPL-X parameters to avoid great motion differences between frames:
$\mathcal{L}_{\text{smooth}}=\lambda_{\text{smooth}}\sum\limits_{i}\sum\limits_{j\in\{1,2\}}(a_{\theta_{j,i}})^{2}+(a_{T_{j,i}})^{2}+(a_{O_{j,i}})^{2}.$ (19)
To leverage contact constraints, we attract human hands to the corresponding contact region on target. We select the positions of 20 fingertips of the two persons in the $i$-th frame as $\mathcal{H}_{i}=\{\bar{P}_{\text{tip,i}}\}_{\text{tip}\in[1,20]}$, where $\bar{P}$ are tip positions in the object’s coordinate system. The contact vertices on the target from object-centric contact retargeting are defined as $\mathcal{C}=\{\bar{P}^{\prime}_{\text{tip}}\}_{\text{tip}\in[1,20]}$. We minimize the Chamfer Distance ($CD$) between $\mathcal{H}_{i}$ and $\mathcal{C}$ to obtain contact consistency:
$\mathcal{L}_{c}=\lambda_{c}\sum\limits_{i}CD(\mathcal{H}_{i},\mathcal{C}).$ (20)
The total human motion retargeting problem is:
$\theta_{1,2},T_{1,2},O_{1,2}\longleftarrow\mathop{\operatorname{argmin}}\limits_{\theta_{1,2},T_{1,2},O_{1,2}}(\mathcal{L}_{j}+\mathcal{L}_{c}+\mathcal{L}_{\text{spat}}+\mathcal{L}_{\text{smooth}}),$ (21)
In practice, we run 1,000 and 1,500 iterations respectively for object motion retargeting and human motion retargeting. The whole pipeline is implemented in PyTorch with Adam solver. The learning rate is 0.01. In object motion retargeting, $\lambda_{f}$ for rotation is 500, for translation is 0.005, $\lambda_{\text{spat}}=0.01$, $\lambda_{\text{smooth}}=1$. In human motion retargeting, $\lambda_{\text{sr}}=0.1$, $\lambda_{\text{wr}}=0.003$, $\lambda_{c}=1,000$, $\lambda_{\text{spat}}=0.01$, and $\lambda_{\text{smooth}}=1$.
C.4 Human-centric contact selection
The pairwise training dataset utilized for the human pose discriminator training comprises 636,424 pairs of data. Each pair encompasses a positive human pose $S_{\text{pos}}\in\mathbb{R}^{21\times 3}$ and a negative human pose $S_{\text{neg}}\in\mathbb{R}^{21\times 3}$. The positive human pose is sampled from the CORE4D-Real. Conversely, the negative human pose is derived from the corresponding positive sample by introducing noise to its object pose, subsequently employing the original contact information to perform contact-guided interaction retargeting. The discriminator is trained by:
$\mathcal{L}_{\text{ranking}}=-\log(\sigma(R_{\text{pos}}-R_{\text{neg}}-m(S_{\text{pos}},S_{\text{neg}}))),$ (22)
iterating 1,000 epochs by the Adam solver with a learning rate 2e-4.
Specifically, the noise $\Delta(\alpha,\beta,\gamma,x,y,z)$ incorporates both rotational and translational components. The rotational noise $\Delta(\alpha,\beta,\gamma)$ ranges from 20 to 60 degrees, while the translational noise $\Delta(x,y,z)$ falls within the range of 0.2 to 0.5 meters. The margin is computed by:
$m(S_{\text{pos}},S_{\text{neg}})=(|\alpha|+|\beta|+|\gamma|)/10+(|x|+|y|+|z|)*10.$ (23)
During the contact constraint update process, a penetration filtering step is performed. For each frame, the penetration volume between the human and object is calculated. If the penetration volume exceeds $10^{-4}$ cubic meters, it is considered a penetration case. If more than 2.5% of frames within an interaction candidate exhibit penetration, the entire candidate is discarded. Among the remaining candidates, the one with the highest score from the human pose discriminator is selected to proceed with the contact constraint update.
Appendix D Dataset Statistics and Visualization
Set ##\##Object ##\##Sequence
Chair Desk Box Board Barrel Stick Chair Desk Box Board Barrel Stick
Real 5 6 9 5 9 4 157 213 200 128 206 58
Synthetic 418 408 376 589 602 596 1767 1344 1326 2123 1495 1961
Table 8: Statistics on object in CORE4D.
Figure 6: T-SNE visualization of human poses for different collaboration modes.
D.1 Collaboration Modes
CORE4D encompasses five human-human cooperation modes in collaborative object rearrangement. “Move1” refers to the scenario where two participants simultaneously rearrange objects and both are aware of the target. On the other hand, “move2” represents the scenario where objects are rearranged simultaneously, but only Person 1 knows the target. “Pass” indicates that one participant passes the object to another for relay transportation. “Join” means that Person 2 joins Person 1 in carrying the object during transportation. Lastly, “leave” signifies that Person 2 leaves during the joint transportation with Person 1.
According to the different durations of the two participants’ contact with the object, “move1” and “move2” can be combined into collaborative carrying tasks. “Pass” represents the task of handover and solely moving the object. Incorporating the join task and the leave task, CORE4D totally comprises four different tasks (see Figure 4 in the main paper) based on the interaction between humans and objects. Fig. 10 exemplifies the motions for each task.
As depicted in Fig. 6, distinct characteristics are exhibited by different cooperation modes in high-level movements, thereby offering an innovative standpoint and potential for comprehending and investigating collaborative behaviors.
D.2 Participants
31 participants, encompassing variations in height, weight, and gender, contributed to the capturing of CORE4D-Real.
D.3 Objects
CORE4D-Real has 38 objects while CORE4D-Synthetic has about 3k objects. The objects encompass six categories, namely box, board, barrel, stick, chair, and desk, each exhibiting a rich diversity in surface shape and size. The distribution of object categories is detailed in Table 8. All the objects in CORE4D-Real are shown in Fig. 9. Fig. 8 shows samples from CORE4D-Synthetic and their interpolation process.
D.4 Camera Views
Fig. 7 shows the four allocentric and one egocentric views of our data capturing system.
Figure 7: Visualization of CORE4D camera views.
Figure 8: Visualization of CORE4D-Synthetic objects and interpolation.
Appendix E Details on Data Split
Benefiting from the diverse temporal collaboration patterns from CORE4D-Real and the large data amount of CORE4D-Synthetic, we randomly select a subset of real object models and construct the training set as the combination of their real (T-Real) and synthesized (T-Synthetic) collaboration motion sequences. We formulate two test sets on CORE4D-Real supporting studies of both non-generalization and inner-category generalization. The first test set (S1) consists of interaction performed on the objects that appear in the training set, while the second one (S2) is composed of interaction from novel objects. Detailed data distribution of each object category is shown in Table 9.
Set ##\##Object ##\##Sequence
Chair Desk Box Board Barrel Stick Chair Desk Box Board Barrel Stick
T-Real 3 4 6 3 6 2 93 104 96 51 113 25
T-Synthetic 418 408 376 589 602 596 1767 1344 1326 2123 1495 1961
S1 3 4 6 3 6 2 40 62 45 21 51 6
S2 2 2 3 2 3 2 24 47 59 56 42 27
Table 9: Train-test split on CORE4D.
Appendix F Evaluation Metrics for Benchmarks
F.1 Human-object Motion Forecasting
Evaluation metrics include the human joints position error $J_{e}$, the object translation error $T_{e}$, the object rotation error $R_{e}$, the human-object contact accuracy $C_{\text{acc}}$, and the penetration rate $P_{r}$.
  • We define $J_{e}$ as the average Mean Per Joint Position Error (MPJPE) of the two persons. MPJPE represents the mean per-joint position error of the predicted human joint positions and the ground-truth values. The unit of $J_{e}$ is one millimeter.
  • Translation error ($T_{e}$) and rotation error ($R_{e}$) denote the average L2 difference between the predicted object translation vectors and the ground-truth ones, and the average geodesic difference between the estimated object rotation matrices and the ground-truth ones, respectively. The unit of $T_{e}$ is one millimeter. The unit of $R_{e}$ is one degree.
  • Physical metrics: To assess contact fidelity, we detect contacts on the two hands of the two persons for each frame with an empirically designed distance threshold (5 centimeters). We then examine the contact accuracy ($C_{acc}$), which indicates the average percentage of contact detection errors in the predicted motions. Additionally, we examine the object penetration ratio ($P_{r}$) representing the mean percentage of object vertices inside the human meshes. The units of the two metrics are percentages.
F.2 Interaction Synthesis
Following an existing individual human-object interaction synthesis study [53], the evaluation metrics include the root-relative human joint position error $RR.J_{e}$, the root-relative human vertex position error $RR.V_{e}$, the human-object contact accuracy $C_{\text{acc}}$, and the FID score (FID).
  • $RR.J_{e}$ denotes the average root-relative MPJPE of the two persons. The root-relative MPJPE represents the mean per-joint position error of the predicted human joint positions relative to the human root position and the ground-truth values. The unit of $RR.J_{e}$ is one millimeter.
  • $RR.V_{e}$ denotes the average root-relative Mean Per Vertex Position Error (MPVPE) of the two persons. The root-relative MPVPE represents the mean per-vertex position error of the predicted human vertex positions relative to the human root position and the ground-truth values. The unit of $RR.V_{e}$ is one millimeter.
  • $C_{\text{acc}}$ is the same as that in Section F.1.
  • The Fréchet Inception Distance (FID) quantitatively evaluates the naturalness of synthesized human motions. We first train a feature extractor on CORE4D-Real to encode each human-object-human motion sequence to a 256D feature vector $\bar{f}_{i}$ and acquire the ground-truth human motion feature distribution $\bar{D}$=$\{\bar{f}_{i}\}$. We then replace the motions of the two persons as synthesized ones and obtain another distribution $D$=$\{f_{i}\}$. Eventually, the FID denotes the 2-Wasserstein distance between $\bar{D}$ and $D$. Since CORE4D-Real provides action labels, the feature extractor is supervised-trained by fulfilling the action recognition task. The network structure of the feature extractor is a single-layer Transformer [101].
Appendix G Qualitative Results on Benchmarks
Figure 11 and Figure 12 exemplify generated motions for the human-object motion forecasting task and the interaction synthesis task, respectively. Since the baseline methods do not focus on generating hand poses, we replace hand poses in ground truth with flat hands to facilitate fair comparisons. Despite diverse cooperation modes that can be generated, the baseline methods consistently encompass unsatisfactory performances including unnatural collaboration, inter-penetration, and unnatural contact.
Appendix H Details on the Application of CORE4D-Synthetic
To evaluate the application of CORE4D-Synthetic, we use the lightweight CAHMP [16] to conduct the motion forecasting experiments. Unlike the experiments in section Human-object Motion Forecasting mentioned in the main paper, where 15 frames are predicted, here we predict the human-object motion for the next 10 frames given the previous 10 frames.
H.1 Task Formulation
Given the object’s 3D model and human-object poses in adjacent 10 frames, the task is to predict their subsequent poses in the following 10 frames. The human pose $P_{h}\in\mathbb{R}^{23\times 3}$ represents the joint rotations of the SMPL-X [79] model, while the object pose $P_{o}=\{R_{o}\in\mathbb{R}^{3},T_{o}\in\mathbb{R}^{3}\}$ denotes 3D orientation and 3D translation of the rigid object model.
H.2 Evaluation Metrics
Following existing motion forecasting works [14, 102, 119], we evaluate human joints position error $J_{e}$, object translation error $T_{e}$, object rotation error $R_{e}$. Details of the three metrics can be found in Section F.1.
H.3 Results
Comparing the 1K real dataset with the 0.1K real dataset supplemented with synthetic data generated through retargeting, we observed that the quality of the synthetic data is comparable to the real data. Additionally, due to the increased diversity of objects and enriched spatial relations between humans and objects in the synthetic data, it exhibits better generalization performance in object motion forecasting.
Comparing the evaluation results of the 1K real dataset with the results obtained by augmenting it with additional 4K synthetic data, we observed a significant performance gain from the synthetic data. This demonstrates that the inclusion of synthetic data enhances the value of our dataset and better supports downstream tasks.
Figure 9: Visualization of CORE4D-Real objects.
Figure 10: Visualization of CORE4D object rearrangement tasks.
Figure 11: Qualitative results of human-object motion forecasting. Grey meshes are from the task inputs.
Figure 12: Qualitative results of interaction synthesis.
Appendix I Details on Humanoid Skill Learning using CORE4D
As introduced in Section 5.2 in the main paper, we use CORE4D’s human-box interaction data to facilitate humanoid skill learning for box lifting. This section presents details on this application, including the simulation environment configuration (Section I.1), task formulation and evaluation (Section I.2), adapting human interaction data to humanoid (Section I.3I.4), benchmark method designs (Section I.5), and experiments (Section I.6).
I.1 Environment Setup
We use the popular Unitree H1 humanoid robot [100] in Isaac Gym [64] simulation environment. The H1 humanoid has 19 revolute joints with fixed limits on motion ranges. The interaction scene contains the humanoid, a box weighing 0.5kg initially posed on the floor, and a third-person-view camera providing visual signals for the skill policy.
Data clipping: The original interactions in CORE4D are collaborations of two persons. In this application, we regard each of them as two individual human-object interaction motions and obtain the box-lifting clips automatically by measuring hand-object distances and the object’s height. We discard the motion if the individual is not touching the object during its lifting process.
Train-test split: As a result, we acquire 890 individual human-box lifting data clips covering 22 boxes augmented from three real-world ones. We select 16 boxes (504 clips) as the training set and the remains as the test set. The training set is used only to train the skill policy (Section I.5) and its preceding motion tracker (Section I.4), while the test set is used only to evaluate the skill policy to assess its generalizability to unseen box shapes.
Observations: The real-time input of the humanoid skill policy is =$\mathcal{I}=\{s_{\text{proprio}},r_{\text{root}},t_{\text{root}},\mathcal{C},\mathcal{D}\}$, where $s_{\text{proprio}}\in\mathbb{R}^{19\times 2}$ is humanoid’s proprioception, $r_{\text{root}}\in\mathbb{R}^{3}$ and $t_{\text{root}}\in\mathbb{R}^{3}$ are humanoid’s root orientation and position in the world coordinate system respectively, and $\mathcal{C}\in\mathbb{R}^{360\times 480\times 3}$ and $\mathcal{D}\in\mathbb{R}^{360\times 480}$ are the color and depth image captured by the camera. Specifically, $s_{\text{proprio}}$ is composed of the angles and velocities of each joint.
Action: Given the real-time input $\mathcal{I}$, the skill policy needs to generate an action $\mathcal{A}$ and use it to actuate the humanoid in the simulation environment. The action $\mathcal{A}\in\mathbb{R}^{19}$ is defined as 19 DoF joint angles, which are transferred to joint torques via a PD controller with pre-defined proportional-derivative gains [26]. The simulation environment uses the computed joint torques to actuate the humanoid.
I.2 Task Formulation and Evaluation Metric
Given an unseen box in the scene and a starting position, the humanoid is required to adjust its pose, touch the box, and finally lift it larger than 20 centimeters. The evaluation metric (SR) is the task success rate defined as whether the box reaches 20 centimeters higher than its initial position.
I.3 Retargeting Interactions from CORE4D onto H1 Humanoid Robot
Following existing humanoid skill learning advances [35, 36, 94], we use an optimization strategy to solve the retargeting problem. The optimization is finding the optimal sequence of $\{s_{\text{proprio}},r_{\text{root}},t_{\text{root}}\}$ that can minimize differences between motions of the human $A$ and the humanoid $B$ on positions of $K$ paired joints in the world coordinate system $L_{p}=\sum\limits_{k=1}^{K}\|P_{a_{k}}-P_{b_{k}}\|^{2}$, where $P\in\mathbb{R}^{T\times 3}$ denotes the position sequence of a joint, and $T$ is the frame number. $<a_{k},b_{k}>$ is a pre-defined pair of joints with similar semantics, where $a_{k}$ is from the human skeleton, and $b_{k}$ is from the humanoid skeleton.
We design a temporal loss $L_{t}=\sum\limits_{t=2}^{T}\sum\limits_{i=1}^{19}(\|s_{\text{proprio},t,i}-s_{\text{proprio},t-1,i}\|^{2}+\|r_{\text{root},t,i}-r_{\text{root},t-1,i}\|^{2}+\|t_{\text{root},t,i}-t_{\text{root},t-1,i}\|^{2})$ to further improve motion smoothness. The overall optimization target is $L_{p}+L_{t}$.
As a result, this retargeting method converts human-box interaction motions to humanoid-box interaction animations. The animations are not physically realistic, and we use the upcoming tracking method to obtain physically realistic ones following HumanPlus [26].
I.4 Tracking Humanoid-box Interaction Animations
Given a humanoid-box interaction animation, the motion tracking methods [82, 62, 96] control the humanoid in the simulation environment and generate the motion that mostly closely resembles the animation. The generated motions are both physically realistic and able to fulfill the task via mimicking animations, which are utilized as skill demonstrations in the following skill policy learning.
We select HST [26] as our tracker. The official implementation of the HST [26] tracker fails to lift the box, and we design an improved HST that can successfully track the animations and control the humanoid to lift the box physically. We describe the key designs of our tracker below.
Input: In each frame, the tracker inputs the states of the real-time humanoid $S_{h}$, the humanoid animation target $\hat{S_{h}}$, the real-time box $S_{b}$, and the box target $\hat{S_{b}}$. $S_{h}=\{r_{\text{root}},\dot{r}_{\text{root}},t_{\text{root}},\dot{t}_{\text{root}},J,\dot{J},R,\dot{R},T,\dot{T}\}$, where $r_{\text{root}}\in\mathbb{R}^{3}$, $\dot{r}_{\text{root}}\in\mathbb{R}^{3}$, $t_{\text{root}}\in\mathbb{R}^{3}$, $\dot{t}_{\text{root}}\in\mathbb{R}^{3}$, $J\in\mathbb{R}^{19}$, $\dot{J}\in\mathbb{R}^{19}$, $R\in\mathbb{R}^{(19+1)\times 3}$, $\dot{R}\in\mathbb{R}^{(19+1)\times 3}$, $T\in\mathbb{R}^{(19+1)\times 3}$, $\dot{T}\in\mathbb{R}^{(19+1)\times 3}$ denotes root’s global orientation, root’s global angular velocity, root’s global position, root’s global linear velocity, joint angle, joint velocity, global positions of the root and joints, global linear velocities of the root and joints, global orientations of the root and joints, and global angular velocities of the root and joints. $\hat{S_{h}}$ is formulated the same as $S_{h}$. $S_{b}=\{r_{b},\hat{r}_{b},t_{b},\hat{t}_{b}\}$, where $r_{b}$, $\hat{r}_{b}$, $t_{b}$, $\hat{t}_{b}$ denotes box global orientation, box global angular velocity, box global position, and box global linear velocity, respectively. $\hat{S_{b}}$ shares the same definition with $S_{b}$.
Output: In each frame, the tracker outputs a 19 DoF action vector representing target joint angles. The joint torques are computed via a low-level PD controller and are fed into the simulation environment to actuate the humanoid.
Reward function: The tracker is trained via a reinforcement learning (RL) method PPO [87]. The reward design is the most crucial part of the method’s performance. Using original rewards from HumanPlus [26] cannot lift the box successfully due to inaccurate tracked hand positions. To handle this issue, we draw inspiration from PhysHOI [105] and use a multiplication of humanoid reward $\mathcal{R}_{h}$, box reward $\mathcal{R}_{b}$, and humanoid-box interaction reward $\mathcal{R}_{i}$ as the overall reward $\mathcal{R}_{\text{overall}}$: $\mathcal{R}_{\text{overall}}=\mathcal{R}_{h}\times\mathcal{R}_{b}\times\mathcal{R}_{i}$.
  • $\mathcal{R}_{h}=0.5\exp(-5\|T-\hat{T}\|_{1})+5\exp(-10\|t_{\text{root}}-\hat{t}_{\text{root}}\|_{1})+10\exp(-10\|J_{l}-\hat{J}_{l}\|_{1})+10\exp(-10\|J_{r}-\hat{J}_{r}\|_{1})$, where $l$/$r$ denotes the left/right hand.
  • $\mathcal{R}_{b}=\exp(-10\|t_{b}-\hat{t}_{b}\|_{1})$.
  • $\mathcal{R}_{i}=\exp(-10\|P_{l,h\rightarrow b}-\hat{P}_{l,h\rightarrow b}\|_{1}-10\|P_{r,h\rightarrow b}-\hat{P}_{r,h\rightarrow b}\|_{1})$, where $P_{h\rightarrow b}$ and $\hat{P}_{h\rightarrow b}$ denotes the humanoid joint position in the box’s coordinate system, and $l$/$r$ denotes the left/right hand.
Training strategies: We adopt an early-termination strategy from DeepMimic [82] that terminates the rollout when the humanoid is 0.5 meters away from its target or its root height is below 0.5 meters.
I.5 Reinforcement Learning and Imitation Learning Method Designs
Combining the retargeting method (Section I.3) with the improved HST tracker (Section I.4), we transfer CORE4D’s data to physically realistic humanoid box-lifting demonstrations. The final step is to train a skill policy that mimics the demonstrations and can lift unseen boxes in test time. We select two vision-based imitation learning (IL) methods, HIT [26] and ACT [143], and use their official implementations.
To examine the value of the demonstrations, we compare the two IL methods with a commonly used state-based RL algorithm PPO [87]. The PPO is implemented with the code from HumanPlus [26], with a change on the reward design $\mathcal{R}$: $\mathcal{R}=\mathcal{R}_{b}+\mathcal{R}_{\textbf{success}}+\mathcal{R}_{i}+\mathcal{R_{\textbf{alive}}}$, where:
  • $\mathcal{R}_{b}=\exp(\|t_{b}-\hat{t}_{b}\|_{2}^{2})$, where $\hat{t}_{b}$ is the pre-defined target box center position for the task.
  • $\mathcal{R}_{\textbf{success}}=[\mathcal{R}_{d}<0.01]$ encouraging achieving the task.
  • $\mathcal{R}_{i}=-0.1(\|P_{l,h\rightarrow b}\|_{2}^{2}+\|P_{r,h\rightarrow b}\|_{2}^{2})$, where $P_{h\rightarrow b}$ and $\hat{P}_{h\rightarrow b}$ denotes the humanoid joint position in the box’s coordinate system, and $l$/$r$ denotes the left/right hand. This reward encourages humanoid hands to explore near the object.
  • $\mathcal{R_{\textbf{alive}}}=0.1$ encouraging the humanoid being alive.
I.6 Experiments
[…]
Appendix J CORE4D-Real Data Capturing Instructions
Target. We divide a $4m\times 5m$ field into 20 squares and number them, and place colored labels as markers along the perimeter of the field. The following language instructs participants: "Please collaboratively move the object to the target square. You can choose any path and orientation of the object as you like. It is not necessary to be overly precise with the final position - a rough placement is fine. Do not make unnatural motions just to achieve an exact position. Do not use verbal communication with each other.". As for the settings when only one participant knows the target, the target square number is written on a piece of paper and shown to the participant who knows the target. And additional instructions are given as: "If you know the target, do not use language or direct body language to inform the other party (such as pointing out the location). If you do not know the target, please assist the other participant in completing the transportation.".
Collaboration Mode. The instructions are given as follows to indicate different Collaboration Modes for the participants. For Collaborate mode: "Based on the target, please cooperatively transport the object, or upright any overturned tables, chairs, etc. Both participants should be in contact with the object throughout the process.". For Handover mode: "Please decide the handover point yourselves, then have one person hand the object to the other, completing the object transfer in relay.". For Leave and Join modes: "One person will transport the object throughout, while the other leaves or joins to help at a time point not disclosed to the collaborator.".
Obstacle. The instructions are given as follows to guide the participants in tackling obstacles: "There are a varying number of obstacles on the field. If they get in your way, please decide on your own how to solve it using some common everyday operations. If the obstacles occupy the destination, please place the object near the destination.".